Find the slope of the line determined by the following equation:
x= 5-y/4
Please show steps to solving.
Thank you.

Answer by Greenfinch(383)

x = 5 - y/4 multiply by 4 to get
4x = 20 - y rearrange
y = 20 - 4x
Gradient is therefore -4. When x = 0, y = 20 so
y intercept +20